Session: Nov 2003 (session 3)
Wavelength: 6 cm
To give feedback on this experiment, click
Torun (K.Borkowski at Wed Nov 5 07:35:25 2003) Failure : RECO
Our P&G recorder is out of order; see the feedback for N03C1.
Shanghai (Xinyong Huang at Wed Nov 5 08:51:36 2003) Success :
The data is invalid from UT:04:42:20 to UT:04:47:26, because of the antenna with trouble.
Urumqi (yang wenjun at Wed Nov 5 14:00:23 2003) Minor failures :
we used the uncooling-receiver.
Medicina (mauro at Wed Nov 5 15:21:49 2003) Success :
All scan recorded.
Effelsberg (A. Kraus at Wed Nov 5 16:35:07 2003) Success :
Onsala (Michael Lindqvist at Thu Nov 6 09:57:06 2003) Minor failures :
Main problem with IF2 solved just after the experiment started,
missed first scans. Still TSYS for IF2 is higher than normal.
Jodrell_Bank (Alastair Gunn at Fri Nov 7 02:41:31 2003) Success :
No known problems.
Hartebeesthoek (Jonathan Quick at Fri Nov 7 21:42:33 2003) Success :
Westerbork (tony foley at Mon Nov 10 09:49:45 2003) Success :
no known problems
Noto (P. Cassaro at Mon Nov 10 16:12:40 2003) Success :